OpenStudy (anonymous):

Solve, but please show work: x^2/12 - 2x/3 - 4=0

OpenStudy (anonymous):

\[\Large \frac{x^2}{12}-\frac{2x}{3}-4=0 \] do you mean this? :)

myininaya (myininaya):

\[\frac{1}{12}x^2-\frac{2}{3}x-4=0\]

OpenStudy (anonymous):

It just says math processing error

OpenStudy (anonymous):

That's quadratic equation... you have \[\Large a=\frac{1}{12} \quad b=-\frac23 \quad c=-4\] use this formula... \[\Large x_{1/2}=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] just substitute and grab calculator (if you need it ! :P) LOL ;)

OpenStudy (anonymous):

Okay thanks(:

myininaya (myininaya):

@dylan82 are you still having a problem viewing the math stuff?

Directrix (directrix):

x^2/12 - 2x/3 - 4=0 Multiply each member of the equation by 12 to clear fractions. x^12 - 8x - 48 = 0 This factors to: (x - 12 ( x + 4) = 0 By the Zero Product Property x - 12 = 0 or x + 4 = 0 x = 12 or x = -4

OpenStudy (anonymous):

@myininaya Not any more